OpenStudy (anonymous):

Sarah’s dad tells her she can have the greater of the following two amounts for her allowance (in dollars): 8 minus 3 times 2 8 less than 3 times 2 What is Sarah’s allowance, in dollars?

hero (hero):

8 - (3*2) = 8 - 6 = 2

jimthompson5910 (jim_thompson5910):

8 minus 3 times 2 = 8 - 3*2 = 8-6 = 2 and 8 less than 3 times 2 = 3*2 -8 = 6-8 = -2 The greater of the two is the first one, so her allowance is $2

OpenStudy (anonymous):

stingy

hero (hero):

lol

OpenStudy (anonymous):

thankx ya'll